OpenStudy (anonymous):

Solve using variation of parameters \[y''-y'=e^x\] \[r^2-r=0\] \[r_1=r_2=1\] \[y_c=c_1e^x+c_2xe^x\] \[W(y_1,y_2)=0\] Is that possible?

OpenStudy (anonymous):

No I did that wrong, one moment

OpenStudy (anonymous):

no it equals \[e^{2x}\]

OpenStudy (anonymous):

\[y_p(x)=-e^xx^2+x^2e^x\]

OpenStudy (turingtest):

wrong complimentary

OpenStudy (turingtest):

\[r^2-r=r(r-1)=0\implies r=\{?,?\}\]

OpenStudy (anonymous):

\[r^2-r=0\]

OpenStudy (anonymous):

r=0 and r=1

OpenStudy (turingtest):

so the complimentary is...?

OpenStudy (anonymous):

\[y_c=c_1+c_2e^x\]

OpenStudy (turingtest):

right, now I get a bit dicey.... since you have e^x on the RHS as well I feel you may need to multiply this by x to keep it linearly independent

OpenStudy (anonymous):

sure \[y_c=c_1+c_2xe^x\]

OpenStudy (turingtest):

\[y_c=c_1x+c_2xe^x\]

OpenStudy (anonymous):

why on both c_1 and c_2, how did come to that conclusion?

OpenStudy (turingtest):

or do you do it to the particular instead? hm... well normally you do it to the whole RHS particular, but since you need the complimentary for the particular in VP I'm not sure...

OpenStudy (anonymous):

I'm not sure... DunnBros is about to kick me out, they close at 11pm

OpenStudy (anonymous):

I'll be back on in like 15 mins when I get home

OpenStudy (turingtest):

fair enough, I'm about to call it a night as well see ya, I will investigate this :)

OpenStudy (anonymous):

Make that 7 mins, ha!

OpenStudy (anonymous):

\[W(y_1,y_2)=x^2e^x\]

OpenStudy (anonymous):

\[y_1=x\] \[y_2=xe^x\]

OpenStudy (anonymous):

\[y_p(x)=-x\int \frac{e^x}xdx+xe^x\int x^{-1}dx\]

OpenStudy (wasiqss):

well ik it

OpenStudy (wasiqss):

my twin :D

OpenStudy (anonymous):

hi twin

OpenStudy (turingtest):

that first integral is not going to be so nice I don't think.... you know who knows exactly what we're missing @lalaly DE help please!!!!!

OpenStudy (wasiqss):

well i can do it

OpenStudy (wasiqss):

tell me equation for which i have to solve\y complimentary

OpenStudy (anonymous):

y''-y'=e^x

OpenStudy (turingtest):

we need variation of parameters @wasiqss

OpenStudy (wasiqss):

yehh just tell me equation i ll take out solution. as y complimentary is solved i ll solve for y particular

OpenStudy (turingtest):

Our problem is that the solutions to the complimentary are not independent from the particular, which you can tell since we got W=0 on our first try

OpenStudy (turingtest):

ok @wasiqss let's see what you got :)

OpenStudy (wasiqss):

tell me equation man completely cause this seems messy

OpenStudy (turingtest):

y''-y'=e^x

OpenStudy (wasiqss):

easy one lol

OpenStudy (turingtest):

then do it, by all mean!

OpenStudy (wasiqss):

y complimentary is y= c1 +c2e^-x

OpenStudy (turingtest):

^yes, answer not important, we need the process...

OpenStudy (turingtest):

complimentary is y=c1+c2e^x

OpenStudy (anonymous):

i disagree with the above answer, shouldn't it be e^x, yeah what turing has

OpenStudy (wasiqss):

yes cause roots are zero and -1

OpenStudy (turingtest):

no they are not

OpenStudy (turingtest):

r^2-r=0 r(r-1)=0 r={0,1}

OpenStudy (anonymous):

do you think my W is right? x^2 e^x

OpenStudy (turingtest):

that I doube @MathSofiya because we have e^x on both left and right our solutions are linearly dependet (that means we can't solve it yet) so maybe if we multiply the g(x) by x ? not sure, but worth a try

OpenStudy (turingtest):

with W=e^2e^x I don't think that first integral is closed, check the wolf, I have not yet

OpenStudy (turingtest):

w=x^2e^x *

OpenStudy (anonymous):

do we agree that \[y_c=c_1x+c_2xe^x\] or should it be \[y_c=c_1+c_2e^x\]

OpenStudy (anonymous):

yeah that's what I had

OpenStudy (turingtest):

that I am not sure, but based on our attempt to use the second one failing I would now try the first (we are discovering this together right now :)

OpenStudy (turingtest):

however if we are to try the first we need to multiply g(x) by x (I'm thinking....)

OpenStudy (turingtest):

so let's try that

OpenStudy (turingtest):

otherwise we get W=0 which is a failure

OpenStudy (turingtest):

@lalaly I know you can help us, please where are you?

OpenStudy (anonymous):

\[y_c=c_1x+c_2xe^x\] \[\left|\begin{matrix}x&xe^x\\1&e^x(x-1)\end{matrix}\right|=x^2e^x\]

OpenStudy (anonymous):

can you explain the g(x) concept again and why it being e^x would be a problem because that cancels something out?...

OpenStudy (turingtest):

no I think that is wrong because that leads to the particular having\[\int\frac{e^x}dxx\]

OpenStudy (turingtest):

*\[\int\frac{e^x}xdx\]

OpenStudy (turingtest):

here is the problem... have you taken linear algebra?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@lalaly is here

OpenStudy (turingtest):

okay. then we have a set of vectors (in this case equations) if their determinant is zero we know they are linearly dependent

OpenStudy (wasiqss):

now listen

OpenStudy (wasiqss):

for y particular

OpenStudy (turingtest):

that is a bad thing in DE's we need all of our solutions to be linearly independent, so the determinant of the solutions, (the wronskian) must not be zero, otherwise our solutions are linearly dependent

OpenStudy (wasiqss):

is it xe^x

OpenStudy (wasiqss):

so as the denominator becomes zero

OpenStudy (turingtest):

and would you care to explain how/why ?

OpenStudy (wasiqss):

we multiply by x

OpenStudy (wasiqss):

cofficeint of e^1x

OpenStudy (wasiqss):

=1

OpenStudy (wasiqss):

[xe^]/[(-1)^2-1)]

OpenStudy (wasiqss):

HENCE we get zero in denominator

OpenStudy (turingtest):

anyway... to keep the particular linearly independent from the complimentary we often need to multiply by x, so that is what I am thinking must happen somehow @lalaly will confirm if I am making any sense or not now

OpenStudy (wasiqss):

now we need to apply binomeal expansion

OpenStudy (turingtest):

@wasiqss sorry to be so direct, but no we do not need any binomial expansion here please let's just have a listen to @lalaly

OpenStudy (wasiqss):

arghh i wish i was good at making you understand it

OpenStudy (wasiqss):

we need it!

OpenStudy (turingtest):

you cannot have 0 in the denom, that shows you that you have already made a mistake !!!!

OpenStudy (wasiqss):

well to counter this zero!! we need to multiply x

OpenStudy (turingtest):

...as I was saying...

OpenStudy (wasiqss):

plz no offense but plz go revise DE s

OpenStudy (turingtest):

._.

OpenStudy (wasiqss):

you know what .. we can approach a DE in differnt ways and this one i will certainly approach by binomeal.. cause we studied this way for this case

OpenStudy (turingtest):

yes, but I think it is way more work than is necessary now I just wanna hear @lalaly

OpenStudy (wasiqss):

kay

OpenStudy (turingtest):

oh dear, we seem to have lost her I am going to try this op paper I guess

OpenStudy (wasiqss):

hahaha lana dear

OpenStudy (wasiqss):

\[[1+(D(D+1)-1)^-1[x^2e^x\]

OpenStudy (wasiqss):

we have to expand binomeally [1+(D(D+1)−1)−1

OpenStudy (wasiqss):

well i have a solution on page i wonder how to give you that

OpenStudy (anonymous):

take a picture with your cellphone and attach it

OpenStudy (anonymous):

or use \[\LaTeX\]

OpenStudy (wasiqss):

good idea twin can i do it tomorow because i would need to upload it too?

OpenStudy (turingtest):

okay got it :)

OpenStudy (wasiqss):

latex is something not my type